OpenStudy (anonymous):

In the below system, solve for y in the first equation. x + 3y = 6 2x - y = 10

OpenStudy (kainui):

Solve for y in the first equation? The y is the same in both equations.

OpenStudy (anonymous):

x + 2 x + 6 -x + 2 x + 2 those are my choices

OpenStudy (texaschic101):

solve for y in the first equation......ok... x + 3y = 6 -- subtract x from both sides 3y = -x + 6 -- divide by 3 y = -1/3x + 2

OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -3 2x - 6y = -6

OpenStudy (texaschic101):

5x - 3y = -3 -->(-2) 2x - 6y = -6 -------------- -10x + 6y = 6 (result of multiplying by -2) 2x - 6y = -6 -------------add -8x = 0 x = 0 now sub 0 in for x in either of the equations to find y 5x - 3y = -3 5(0) - 3y = -3 -3y = -3 y = 1 check.. 2x - 6y = -6 2(0) - 6(1) = -6 -6 = -6 (correct) x = 0 and y = 1

OpenStudy (anonymous):

would you do the same here? 4x - y = 3 7x - 9y = -2

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

this one is a bit easier because it can be easily solved with substitution. 4x - y = 3 -y = -4x + 3 -- multiply by -1 to make y positive y = 4x - 3 now just sub 4x - 3 in for y in the other equation to solve for x. Understand ?